in the absence of air resistance, a projectile is launched from and returns to ground level.It has a range of 24.5 m. Suppose the launch is double and the projectile is fired at the same angle above the ground. What is the new range?


1
Expert's answer
2021-02-07T19:19:08-0500
"\frac{R_2}{R_1}=\frac{v_2^2}{v_1^2}\\\frac{R_2}{24.5}=2^2\\R_2=98\ m"